### v2

@@ 2.1.3 How do digital twins compare to benchmarks that require less information?
This finding echoes [18], who argue that LLMs are steerable through customized prompting. In our case, adding digital twin information steers responses in ways consistent with observed variation across human participants (as observed by the improved correlation). However, as [18] emphasize, steerability does not guarantee alignment with human responses. In our results, this manifests as modest or nonexistent gains in individual-level accuracy. To illustrate the distinction between improving correlation and improving individual-level accuracy, we select one outcome (Lack of Control) from one sub-study (Substudy 2, Affective Primes), for which the distinction between individual-level accuracy and correlation is particularly sharp. Figure 3 reports scatter plots of the predictions from synthetic personas created using demographics only vs. human responses (left panel), and predictions from digital twins created using full personas vs. human responses (right panel). These plots clearly illustrate how the correlation is much improved when full personas are used ($r_{demo}=0.105$ vs. $r_{full}=.555$). Yet, individual-level accuracy is virtually unchanged ($accuracy_{demo}=0.892$ vs. $accuracy_{full}=0.907$). Note that individual-level accuracy should not be confused with other aggregate measures of accuracy, such as the accuracy of the average difference between groups of participants (e.g., average treatment effect). Such measures may be impacted more strongly by improvements in correlation.
@@
Figure 3: Human Responses vs. Digital Twin Responses based on Demographics Only (Left) and Human Responses vs. Digital Twin Responses based on Full Persona (Right), for One Particular Outcome. 45° line included. This example illustrates how correlation may be much improved when full personas are used, without significant change to individual-level accuracy.
2.1.4 How do digital twins compare to benchmarks that require more information?
In order to further contextualize the correlation achieved by digital twins, we also compare digital twins to a traditional machine learning approach. Unlike digital twins that are able to predict answers to completely new questions, traditional machine learning models require collecting responses from a subset of the sample, training a model that connects these responses to a set of personal characteristics, and then predicting out-of-sample responses for the rest of the sample (given their personal characteristics). We consider a case in which an XGBoost model is trained separately for each outcome based on the value of that outcome for a sub-sample of participants and their full persona information (i.e., same information as digital twins). [^9] We focus on the 106 outcomes for which sample size is at least 650, and vary the size of the training sample from 50 to 650. We also set temperature to 0 for digital twins, as Section 2.4 reveals this improves their performance. See SI for details. We find that in order for such a traditional machine learning approach to match the out-of-sample predictive correlation achieved by digital twins, one would need to collect each outcome from (i.e., run the actual study on a training sample of) approximately 180 participants on average. Moreover, we find that even as the size of the training sample reaches 650, the predictive correlation achieved by XGBoost remains below 0.29. We also consider an XGBoost model trained on demographic variables only. Such a model never reaches the performance of digital twins: even when using up to 650 participants as training sample, correlation does not exceed 0.18. This analysis suggests that although the correlation achieved by digital twins is modest in absolute terms, it is higher than what could be obtained by traditional machine learning approaches that require additional data. This also suggests that given the type of data we collected, it may be unrealistic to expect correlations higher than 0.3. Future research might further improve digital twin performance by uncovering better types and sources of training data for specific predictive tasks.
Figure 4: Comparison with Traditional Machine Learning Method. We compare digital twins (blue line) to an XGBoost model (orange line) trained for each outcome using additional data not needed or used by digital twins: responses for that outcome from a training subset of participants. In order for such a traditional machine learning approach to match the averagea out-of-sample predictive correlation achieved by digital twins, one would need to collect each outcome from approximately 180 participants. As the size of the training sample reaches 650, the predictive correlation achieved by XGBoost does not exceed 0.29. An XGBoost model trained only on demographic variables (green line) never reaches the performance of digital twins, even when using up to 650 participants as training sample.
2.2 Digital Twin Performance Across Domains
2.2.1 Meta-analysis
We explore patterns in performance across domains using a meta-analytic approach. We conduct separate regressions with each performance metric as the dependent variable. Given that correlation is the performance metric most sensitive to the individual-level information captured by digital twins, we focus here on the regression with the (z-transformed) correlation for each outcome as the dependent variable and report the meta-analysis of the other performance metrics in the SI. [^10] We z-transform the correlation due to better statistical properties, e.g., being approximately normally distributed. Our independent variables include a set of categorical labels that characterize each outcome, related to the following domains: social, preferences/attitudes, cognitive skills/rationality, content evaluation, human-tech interactions. We also include mechanical features that might influence performance (e.g., sample size).
Labels are not mutually exclusive; an outcome may be tagged with multiple labels. All labels are binary, except sample size. The labels are as follows (with examples from our data):
  • Social
    • Social domain (104 outcomes): involves social topics (e.g., likelihood of sharing news content)
    • Social cognition (15 outcomes): involves social cognition, i.e., asking the participant to judge the behavior of others (e.g., judging someone who donates to both political parties)
  • Preferences/attitudes
  • Cognitive skills/rationality
    • Test of rationality (10 outcomes): assesses rational decision-making (e.g., sensitivity to context effect)
  • Content evaluation
    • Valenced (40 outcomes): involves positive vs. negative judgments (e.g., judging the creativity of ideas)
  • Human–tech interactions
  • Mechanical
  • Digital twins performed relatively better in the social domain. The correlation was higher when outcomes related to conflict, pro-social issues, social cognition, or personality. However, the correlation was significantly lower for outcomes where social desirability was salient, suggesting twins are less capable of mimicking human responses in socially-sensitive contexts (a qualitative review suggests twins are more likely to provide socially-desirable responses).
  • The labels related to preferences and attitudes were not significant, except for outcomes related to political preferences: correlations were particularly low for outcomes in the political domain. This is consistent with previous findings such as [18] or [16].
  • Correlation was on average higher in the cognitive domain.
  • Twins did not correlate with humans better or worse on average when evaluating content (i.e., the coefficient associated with content evaluation was not significant in the meta-analysis regressions). However, their responses were less correlated with humans when the evaluation was valenced (positive vs. negative).
  • When questions varied across participants, twin-human correlation dropped, consistent with prior findings that LLMs struggle to treat prompt-level variation as exogenous [9].
  • Twins showed higher correlation with humans on outcomes that use response scales.
Figure 5: Results from Meta-Analysis (Mixed Linear Model with z-transformed Correlation as Dependent Variable). The average correlation across all outcomes is 0.197. Correlation between responses from twins vs. their human counterparts tends to be higher in social domains, except when social desirability is salient. Correlation tends to be higher in the cognitive domain and in domains related to human-technology interactions, but lower in the political domain and when providing valenced evaluations.
2.2.2 Qualitative comparisons
In order to gain more intuition for variations in performance across domains, we complement our meta-analysis at the detailed outcome level (e.g., willingness to share a particular article on social media, likelihood to apply for a particular job) with a more qualitative analysis across 31 pre-registered higher-level comparisons between digital twins and humans (1-4 comparisons per sub-study). These comparisons evaluated whether twins responded similarly to humans on average to a particular question or set of questions (e.g., “How does the knowledge of digital twins regarding fees and surcharges in the marketplace compare to their human counterparts?”), whether the variance in their responses was similar (e.g., “How does the variation in perceptions of fairness regarding fees and surcharges in the marketplace compare between digital twins and their human counterparts?”), whether they replicated known treatment effects (e.g., “Can digital twins demonstrate default effects?”), among others. A complete list of these comparisons and their results are available in the SI; we only highlight a subset here.
Our comparisons suggest that twins tended to express views that were often more “pro-humans” than those expressed by humans. For example, compared to their human counterparts, twins were more likely to believe that people are fair and can be trusted (Substudy 15), that people should take care of themselves (Substudy 15), and they were relatively favorable to people who donate to both political parties (Substudy 16). They were also more favorable to government regulation of fees and surcharges in the marketplace (Substudy 7), and more willing to pay taxes to improve healthcare for all people (Substudy 15). Interestingly, at the same time twins also tended to express views that could be described as “pro-technology.” In particular, twins tended to provide answers that were consistent with a view of technology as a safe tool under the control of humans (e.g., Substudies 6, 9, 14, 19). For example, twins were relatively more accepting of algorithmic hiring compared to their human counterparts (Substudy 9), they perceived online targeting as less intrusive on their privacy (Substudy 14), and they under-reported usage of platforms like Netflix and TikTok (Substudy 19). In cognitive domains, consistent with the hyper-accuracy distortion identified by [1], twins tended to display perfect knowledge on questions with objectively correct answers (Substudies 7,19), and they tended to behave more rationally than their human counterparts (Substudies 17, 19). When evaluating content, twins were not significantly correlated with humans when rating the creativity of ideas, [^11] Interestingly, when told whether the ideas came from humans or from their twins, humans showed AI aversion (rating ideas from twins lower due to their source) but twins showed human aversion (rating ideas from human lower due to their source). See Substudy 10. but they provided more reliable input when evaluating and predicting the content of book chapters (Substudy 8).
2.3 Digital Twin Performance Across Demographic Groups
@@
We next examine whether demographic characteristics of participants predict the accuracy of their digital twins. For each participant, we compute accuracy (across all outcomes) between the participant’s normalized responses and those of their twin. [^12] We switch to accuracy for this analysis as a more natural way to compare the answers between a particular human and their twin across outcomes. We then analyze how this accuracy relates to participants’ demographic characteristics. All demographic variables are dummy-coded, yielding 61 demographic dummy variables coming from the 14 demographic questions. To identify which participant characteristics are most predictive of digital twin performance, we train an XGBoost (Extreme Gradient Boosting) model to predict twin accuracy using these features [5].
@@
To interpret the model’s findings, we generate grouped partial dependence plots (PDPs) for each categorical variable. PDPs visualize the average predicted accuracy for each level of a feature while holding all other variables constant. For example, the PDP for political ideology shows predicted accuracy for participants identifying as "Very Conservative," "Liberal," "Moderate," “Liberal,” and “Very Liberal” on a single chart enabling direct comparison across different demographic groups. Plots for education, income, religious attendance, and political views are presented in Figure 5, while the others are in the SI.
Our findings suggest systematic differences in twin accuracy across demographic groups. Twins tend to be more accurate for participants with higher education levels and higher income, indicating a potential bias in favor of more socioeconomically advantaged individuals which could exacerbate a focus on WEIRD (Western, Educated, Industrialized, Rich, Democratic) populations in social sciences [10]. Accuracy also tends to be higher for participants with moderate political views and religious attendance habits.
@@
Figure 6: Partial Dependence Plots for Understanding Heterogeneity in Digital Twin Performance.Twins tend to be more accurate for participants with higher education levels and higher income, and for participants with moderate political views and religious attendance habits.
2.4 Digital Twin Performance Across Implementation Methods
All results reported so far use GPT4.1 (dated 2025-04-14) with a default temperature of 0.7.131313As an exception, Figure 4 used temperature of 0. We experimented with different temperatures and base LLMs (e.g., GPT-5, Deepseek, Gemini, and a version of GPT4.1 fine-tuned on the Twin-2K-500 dataset). We find the best results overall using GPT4.1 with temperature = 0 (e.g., highest correlation across all implementation options, tied for best on individual-level accuracy). See SI for details.
3 Discussion
@@
Despite the significant interest in leveraging digital twins of humans to simulate surveys and experiments, little is known about how responses from digital twins compare to those of their human counterparts. Further, little is known about the best way to construct digital twins, on the domains of applications in which digital twins are likely to be accurate, or on the profiles of humans whose digital twins are likely to be more accurate. In order to help advance knowledge on digital twins, we ran a mega-study (and make the data publicly available) that: (i) allows matching the answers of digital twins to those of their human counterparts, (ii) covers a wide range of domains of applications, and (iii) involves a relatively large, representative sample of participants.
In a nutshell, we find that while digital twins show promise, they are not fully “ready for prime time” yet. Significant improvements would be needed for twins to be ready for systematic deployment and/or expectations of digital twin performance may need to be appropriately calibrated. We find that the rich individual-level information provided as input to digital twins improves our ability to distinguish between participants and capture heterogeneity, compared to simpler generic personas based on demographic information. The correlation between the responses from twins and their human counterparts is modest (approximately 0.2 on average); however, it is higher than what could be obtained by more traditional machine learning approaches that require additional data, and our analysis suggests it may not be realistic to expect much higher levels of correlation. In our data and with our current pipelines, we also find that the richer individual-level information contained in digital twins does not improve the accuracy of answers at the individual-level, nor does it improve the accuracy of estimates of population-level averages. Finally, we find that while the individual-level information contained in digital twins helps increase the variance of their answers, they are still under-dispersed compared to human answers.
Our meta-analysis further reveals domains in which digital twin responses are currently more strongly correlated with human responses. We find that digital twins are currently better equipped to approximate responses in social domains, except in cases where social desirability is a potential concern. Correlations in the cognitive domain are also higher, but digital twins are more likely to provide objectively-correct answers on questions that require knowledge or cognitive skills. We also find digital twins to be less connected with their human counterparts in the political domain. More research is needed to understand the reliability of twins for content evaluation; our results suggest that twins struggle particularly with valenced evaluations. A more qualitative set of comparisons reveals potential systematic biases, e.g., twins appear to be at the same time “pro-human” and “pro-technology.”
@@
Could human experts predict these results? To assess whether digital twins behaved as human experts would predict, we asked a sample of expert respondents (primarily scholars and managers) to estimate five average treatment effects in our studies representing different paradigms. We asked separately for the average treatment effect for humans and their twins. The SI contains the estimated effects for these five items and significance tests. Overall, the pattern suggested that respondents’ predictions were less accurate for twins than for humans. While the observed average treatment effect among humans was in the 95% confidence interval of predictions made by our participants in all 5 cases, the average treatment effect among twins was outside the 95% confidence interval in 3 cases. This suggests that human experts struggle to predict the behavior of digital twins.
Finally, our analysis suggests that the accuracy of digital twins is uneven across demographic groups, with better alignment for participants who are more educated, higher income, and with moderate political views and religious attendance habits.
A common overarching question that emerges from our results is the role played by the based LLM vs. the individual-level information in shaping the behavior of digital twins. Indeed, a digital twin may be viewed as a combination of a base model with individual-level additional data, i.e., the answers are “shrunk” towards a base model. One may speculate, and test in future research, that several of the results we observe are due to properties of the base LLM being transferred to digital twins. For example, it is likely that the under-dispersion of answers by digital twins is the result of the homogeneity of the base model not being sufficiently offset by the individual-level information contained in digital twins. Similarly, the inability of digital twins to improve individual-level accuracy may be due to systematic errors originating from the base LLM. Some systematic biases revealed by our qualitative comparisons (e.g., pro-human, pro-technology) are also likely to come from the base LLM powering the digital twins. It is similarly quite plausible that the heterogeneity in digital twin performance across participant characteristics is linked to certain groups (e.g., higher education, moderate views) being better represented in the base model’s training data. Of course, a base model is required to enable digital twins to process and generate unstructured data, and to make out-of-sample predictions without additional training data. But future research might find optimal ways to combine the capabilities of one (or potentially an ensemble of) base LLM(s) with rich individual-level information. We hope our data will encourage the development of improved pipelines for digital twins, that will address some of the limitations uncovered in our analysis.
@@
Our findings also raise the question of how we should reasonably think about digital twins. In particular, based on our results it may not be realistic to think about them as “clones” of humans, but rather as hyper-rational, quasi-omniscient versions of humans, with implicit values partly imbued by their base LLM. This may influence which use cases are more promising for future research and practical applications. Perhaps use cases that consider digital twins as well-informed advisors are more promising than use cases that consider digital twins as carbon copies of their human counterparts.

### v4

@@ 1. Stereotyping
It may be counter-intuitive that full personas significantly improve correlation compared to demographics-based personas, without improving accuracy. However, accuracy and correlation are distinct constructs. To illustrate the distinction between improving correlation and improving individual-level accuracy, we select one outcome (Lack of Control) from one sub-study (Substudy 2, Affective Primes), for which the distinction between individual-level accuracy and correlation is particularly sharp. Figure 3 reports scatter plots of the predictions from synthetic personas created using demographics only versus human responses (left panel), and predictions from digital twins created using full personas versus human responses (right panel). These plots clearly illustrate how the correlation is much improved when full personas are used ($r_{demo}=0.105$ versus $r_{full}=.555$). Yet, individual-level accuracy (captured by the average distance of each point to the 45° line) is virtually unchanged ($Accuracy_{demo}=0.892$ versus $Accuracy_{full}=0.907$).
@@
Figure 3: Human Responses versus Digital Twin Responses Based on Demographics Only (Left) and Human Responses versus Digital Twin Responses Based on Full Persona (Right), for One Particular Outcome. The 45° line is included. This scatter plot illustrates how correlation may be improved when full personas are used, without significant change to individual-level accuracy (captured by the average distance of each point to the 45° line).
To illustrate this comparison more concretely, consider two people with similar demographics. One person rates themselves a 2 out of 7 and the other rates themselves a 4 out of 7 on feeling a lack of control. Using demographics only, the twins might report 3 for each, 1 point away from the human’s actual response on average. Using the full persona, the twins might now report 3 and 5 respectively, still 1 point off on average, failing to capture the individual’s subjective experience to a similar extent, but now capturing the relative difference across people. This suggests that personalized twins can sort people more accurately (who feels more or less control) but still fail to capture individuals’ actual cognition and subjective experiences. Note that individual-level accuracy should not be confused with other aggregate measures of accuracy, such as the accuracy of the average difference between groups of participants (e.g., average treatment effect). Such measures may be more strongly impacted by improvements in correlation.
In sum, one way digital twins distort human behavior is by over-relying on demographic information rather than modeling the complexity of individual cognition.
2. Insufficient Individuation
The modest correlation between digital twins and their human counterparts reveals a failure to fully capture heterogeneity across people. One possibility is that digital twins insert the wrong variance– such as relying too heavily on superficial demographic characteristics. Another source of this inaccuracy could be insufficient variation. Indeed, prior work on other forms of synthetic data finds that LLM-generated responses often suffer from excessive homogeneity (?). To examine this, we compare the amount of variation in answers from twins vs. humans. The standard deviation of the twin responses is lower than that of human responses in 154 of 164 cases (93.9%), indicating under-dispersion in twin responses. This difference is statistically significant (p<<0.05) in 146 of those 154 cases.
Adopting a Bayesian framework, one may think of the base LLM as reflecting a prior distribution over how an individual may respond to a particular question. The additional information fed to the LLM to build that person’s digital twin leads to an updated, posterior distribution of answers tailored to that individual. The under-dispersion of digital twin answers relative to their human counterparts suggests that this additional data is able to steer the distribution in an appropriate direction (i.e., the variations across digital twins better mirror the variations across people), but the base model (i.e., the prior) still carries significant weight and influence on the digital twins’ responses. In other words, the answers are overly “shrunk” towards a base model.
To quantify this, we compare the MAD between answers from full-persona twins vs. humans to the MAD between answers from full-persona twins vs. empty-persona twins (which only reflect the base LLM). Ideally, full-persona twins would behave more similarly to humans than to empty-persona twins. We find the opposite: full-persona twins are closer to empty-persona twins than they are to humans ($MAD_{full\ vs.\ empty}=0.175$, $MAD_{full\ vs.\ humans}=0.252$, $p<0.01$). This finding echoes (?), who argue that LLMs are steerable through customized prompting, but that steerability does not guarantee alignment with human responses.
In sum, we find that digital twins also distort human behavior by insufficiently deviating from the base model. One might wonder whether digital twins suffer more from insufficient individuation in certain domains. We explore this in the SI using a meta-analytic approach. For example, we find that digital twins are currently better equipped to capture heterogeneity in responses in social domains, except in cases where social desirability is a potential concern. We also find digital twins to be less correlated with their human counterparts in the political domain, and when generating valenced evaluations.
3. Representation Bias
@@
We have so far demonstrated two distortions that can explain inaccuracy. But is this inaccuracy uniform across the humans that digital twins are attempting to represent? As training data are typically based on corpora that systematically over-represent certain demographics (?), perhaps digital twin inaccuracy is exacerbated for under-represented groups. To explore this possibility, we compute accuracy (across all outcomes) between each participant’s normalized responses and those of their twin. [^9] We use accuracy for this analysis as a more natural way to compare the answers between a particular human and their twin across outcomes. We then analyze how this accuracy relates to participants’ demographic characteristics. All demographic variables are dummy-coded, yielding 61 demographic dummy variables coming from the 14 demographic questions. To identify which participant characteristics are most predictive of digital twin performance, we train an XGBoost (Extreme Gradient Boosting) model to predict twin accuracy using these features (?).
@@
To interpret the model’s findings, we generate grouped partial dependence plots (PDPs) for each categorical variable. PDPs visualize the average predicted accuracy for each level of a feature while holding all other variables constant. For example, the PDP for political ideology shows predicted accuracy for participants identifying as ”Very Conservative,” ”Liberal,” ”Moderate,” “Liberal,” and “Very Liberal” on a single chart, enabling direct comparison across different demographic groups. Plots for education, income, religious attendance, and political views are presented in Figure 4, while the remaining plots are in the SI.
@@
Figure 4: Partial Dependence Plots for Understanding Heterogeneity in Digital Twin Performance. Twins tend to be more accurate for participants with higher education levels and higher income, as well as those with moderate political views and religious attendance habits.
Our findings suggest representation bias, i.e., systematic differences in twin accuracy across demographic groups. Twins tend to be more accurate for participants with higher education levels and higher income, indicating a potential bias in favor of more socioeconomically advantaged individuals, which could exacerbate a focus on WEIRD (Western, Educated, Industrialized, Rich, Democratic) populations in the social sciences (?). Accuracy also tends to be higher for participants with moderate political views and religious attendance habits.
In sum, this differential accuracy constitutes another distortion: digital twins systematically misrepresent certain groups more than others.
4. Ideological Biases
The first three distortions discussed above represent failures to capture individual cognition by over-relying on demographic information, insufficiently adjusting away from the base model, or insufficiently capturing the specific behaviors of certain groups. These can be seen as errors of omission. However, another possible avenue of distortion is the introduction of bias in responses. Our data provide a unique opportunity to contribute to the growing literature on LLM biases by directly comparing opinions expressed by humans to those expressed by an LLM instructed to mimic those humans, across a wide range of substantive areas.
Specifically, in addition to 164 detailed outcomes (e.g., willingness to share a particular article on social media, likelihood of applying for a particular job), we also pre-registered 31 higher-level comparisons between digital twins and humans (1-4 comparisons per sub-study). These comparisons evaluated whether twins responded similarly to humans on average to a particular question or set of questions (e.g., “How does the knowledge of digital twins regarding fees and surcharges in the marketplace compare to that of their human counterparts?”), whether they replicated known treatment effects (e.g., “Can digital twins demonstrate default effects?”), etc. A complete list of these comparisons and their results is available in the SI; we only highlight a subset here, which reveal systematic ways in which the content of digital twin answers differ from that of humans. In particular, our diverse mega-study allowed us to uncover subtle ideological biases embedded in the twins’ responses.
Specifically, we found that twins tended to express views that were often more “pro-human” than those expressed by humans. For example, compared to their human counterparts, twins were more likely to believe that people are fair and can be trusted (Substudy 15), that people should take care of themselves (Substudy 15), and were relatively more favorable to people who donate to both political parties (Substudy 16). They were also more favorable toward government regulation of fees and surcharges in the marketplace (Substudy 7), and more willing to pay taxes to improve healthcare for all people (Substudy 15).
Interestingly, at the same time, twins also tended to express views that could be described as “pro-technology.” In particular, twins tended to provide answers that were consistent with a view of technology as a safe tool under the control of humans (e.g., Substudies 6, 9, 14, 19). For example, twins were relatively more accepting of algorithmic hiring compared to their human counterparts (Substudy 9), perceived online targeting as less intrusive on their privacy (Substudy 14), and under-reported usage of platforms like Netflix and TikTok (Substudy 19). While it is possible to be simultaneously “pro-human” and “pro-technology,” do digital twins show any preference in favor of humans versus technology? When asked to rate the creativity of ideas (Substudy 10), twins showed human aversion (rating ideas from humans lower due to their source).
These biases may come from various inputs into the base model. In particular, they may be reflected in the training corpus of the base model, or they may have been acquired through Reinforcement Learning with Human Feedback (?) or through the base model’s system prompt which provides overarching instructions on how the model should behave. Irrespective of their exact sources, they imply that blindly deploying digital twins might lead to distorted conclusions about human behavior and opinions, which may, sometimes in very subtle ways, push specific ideological agendas.
In sum, we find that digital twins distort human behavior by exerting ideological sway.
5. Hyper-rationality
Another form of bias that may be introduced by digital twins relates to cognitive skills. Specifically, one of the features of LLMs – their vast knowledge base– may be a source of error when it comes to predicting human behavior. Akin to the curse of knowledge in humans (?), digital twins may be unable to “forget” their base knowledge and behave as more informed, rational agents. Indeed, prior work finds that LLMs suffer from hyper-accuracy (?, ?). We first examined this possibility in sub-studies that contained questions with objectively correct answers (Substudies 7,19). Consistent with prior work, we find that twins tended to display perfect knowledge, deviating from human participants.
We next examined this distortion for substudies that contained questions that test rationality. We again find evidence of hyper-rationality distortion: twins were more likely to select normative answers reflective of higher cognitive abilities, e.g., by selecting answers associated with stronger quantitative intuition (Substudy 17) or by being more likely to strategically consume online content in order to influence future recommendations (Substudy 19). Twins also failed to replicate the attraction and compromise effects (Substudy 4), which are indicators of bounded rationality. However, an interesting phenomenon emerges when using a well-known, published paradigm. Here, twins were more likely to show the effect, probably due to the phenomenon of leakage, whereby the answer is directly influenced by information contained in the training corpus. For example, digital twins did not replicate a default effect using a novel paradigm (Substudy 5), but they did display a strong default effect using a classic paradigm (?). Other substudies (e.g., Substudy 1) confirmed the tendency of digital twins to behave consistently with the literature when a clear prediction could be made. This highlights a key contribution of our megastudy: we are the first to systematically test twins across both established and novel paradigms.
In sum, digital twins distort human behavior by imposing a hyper-rational lens.
Discussion
@@
Although many are rushing to simulate surveys and experiments using digital twins of humans, little is known about how responses from digital twins compare to those of their human counterparts. Further, there is little understanding of the best way to construct digital twins, the domains of application in which digital twins are likely to be accurate, or on the profiles of humans whose digital twins are likely to be more accurate. Deploying digital twins without answering these fundamental questions risks misrepresenting human cognition in ways that could undermine both scientific understanding and practical applications. In order to help advance knowledge on digital twins, we ran a mega-study (and make the data publicly available) that: (i) allows matching the answers of digital twins to those of their human counterparts, (ii) covers a wide range of application domains, and (iii) involves a relatively large, representative sample of participants.
We find that although digital twins show promise, they are not yet “ready for primetime.” Digital twins capture individual cognition only minimally better than empty personas. However, twins show slightly more promise when it comes to sorting people. This distinction is consequential for understanding when indiscriminate twin use may mislead decision-makers. For example, although twins may be able to modestly estimate relative receptiveness to a policy (identifying who is more or less likely to respond positively), they are barely better at predicting an individual’s likelihood to agree with a given policy than a base LLM.
@@
Could human experts have predicted these results? To assess whether digital twins behaved as human experts would have predicted, we asked a sample of expert respondents (primarily scholars and managers) to estimate average outcomes in five of our experimental studies representing different paradigms. We asked separately for the average outcome for humans and their twins in a treatment condition (e.g., likelihood of staying enrolled in a green energy plan when enrolled by default), providing the average outcome in the control condition as baseline (e.g., likelihood of actively choosing the green energy plan when enrolled in another plan by default). Overall, the pattern suggested that respondents’ predictions were less accurate for twins than for humans (see SI). Comparing the predicted outcome in the treatment condition to its value in the control condition also yields estimated treatment effects. While the observed average treatment effect among humans was within the 95% confidence interval of predictions made by our participants in all five cases, the average treatment effect among twins was outside the 95% confidence interval in three cases. This suggests that human experts struggle to predict the behavior of digital twins, highlighting the importance of our rigorous examination.
Beyond performance metrics, our analysis highlights five systematic ways in which digital twins distort human responses. First, we find evidence of demographics-based stereotyping. Second, we find that digital twins suffer from insufficient individuation, providing answers that remain under-differentiated and under-dispersed relative to human answers. Third, digital twin predictions suffer from representation bias, as they tend to be more accurate for certain demographic groups of participants with higher education levels, higher income, and moderate political views and religious attendance habits. Fourth, digital twins tend to display ideological biases, for example, they appear to be at the same time “pro-human” and “pro-technology.” Fifth, digital twins tend to display hyper-rationality, with answers reflecting higher rationality and knowledge compared to humans.
These results culminate in the conclusion that the digital twins of today may best be described as funhouse mirrors that systematically distort human behavior. Stakeholders deploying digital twins risk making consequential decisions based on an illusion of personalization rather than genuine individual-level insight. These systems are better characterized as “comparative profiles” imbued with LLM biases, not faithful twins of human cognition. Consequently, the premature deployment of these twins risks smoothing out individual voices, mispredicting ignorance and irrationality, reinforcing stereotypes, and introducing novel biases, such as pro-technology attitudes and AI-favoritism.
Future Directions
How might future research address these current distortions? Recall that leveraging digital twins involves making out-of-distribution predictions using a base LLM augmented with individual-level data (the twin’s “blueprint”).
We compare digital twins to a traditional machine learning approach in order explore the extent to which the modest performance of digital twins (and therefore, future potential improvements) lie in the blueprint data being insufficiently informative vs. the LLM being unable to fully leverage these data for out-of-distribution predictions. Unlike digital twins, which are able to make out-of-distribution predictions on completely new questions, traditional machine learning models require collecting responses from a subset of the sample, training a model that connects a set of personal characteristics to these responses, and then predicting out-of-sample responses for the rest of the sample (given their personal characteristics).
We consider a case in which an XGBoost model is trained separately for each outcome based on the value of that outcome for a sub-sample of participants and their full persona information (i.e., the same information used for full-persona digital twins). [^10] We focus on the 106 outcomes for which sample size was greater than 650, and vary the size of the training sample from 50 to 650. We also set temperature to 0 for digital twins, as this improves their performance (see SI for details). We also consider an XGBoost model trained on demographic variables only. Comparing the performance of a given model with full-persona information vs. demographics-only information sheds light on the intrinsic value of the twins’ blueprint data. Comparing the performance of the XGBoost model to that of digital twins holding the persona information constant sheds light on the base model’s ability to leverage individual-level data for out-of-distribution predictions.
Looking first at correlation (see Figure 5), comparing each model with full personas vs. demographics-only personas reveals that the full-persona blueprint data helps both digital twins and the XGBoost improve performance. However, even as the size of the training sample reaches 650, the predictive correlation achieved by the full-persona XGBoost model remains below 0.29. This suggests that, despite being extensive and based on decades of social science research, the twins’ blueprint data has limited intrinsic potential to capture human behavior in our context. At the same time, we see that full-persona digital twins reach a predictive correlation equivalent to that achieved by the full-persona XGBoost model trained on approximately 180 participants. In the demographics-only case, the number increases to approximately 225 participants. In other words, given a set of individual-level characteristics, the base LLM is able to make out-of-distribution predictions that achieve the same correlation with human answers as out-of-sample predictions made by an XGBoost model trained on approximately 200 human answers (i.e., that require running the actual study on approximately 200 participants).
When it comes to accuracy (Figure 6), the XGBoost model trained on the full data is only marginally better than the XGBoost model trained on demographics only (and similarly for digital twins), suggesting that the twins’ blueprint data is only marginally informative of the idiosyncratic patterns in responses. Moreover, digital twins are only as good as XGBoost trained on approximately 75 human responses on average, suggesting that it is harder for the LLM to leverage these data to make out-of-distribution predictions that would be accurate at the individual level.
Figure 5: Comparison with Traditional Machine Learning Method: Correlation. We compare digital twins (plain lines) to an XGBoost model (dashed lines) trained for each outcome using additional data not needed or used by digital twins: responses for that outcome from a training subset of participants. We consider versions of each model based on full-persona information (darker colors) vs. demographics only (lighter colors). The full-persona information helps both digital twins and the XGBoost improve performance. However, even as the size of the training sample reaches 650, the predictive correlation achieved by the full-persona XGBoost model remains below 0.29. Holding the set of individual-level characteristics constant, the base LLM is able to make out-of-distribution predictions that achieve the same correlation with human answers as out-of-sample predictions made by an XGBoost model trained on approximately 200 human answers.
Figure 6: Comparison with Traditional Machine Learning Method: Accuracy. We compare digital twins (plain lines) to an XGBoost model (dashed lines) trained for each outcome using additional data not needed or used by digital twins: responses for that outcome from a training subset of participants. We consider versions of each model based on full-persona information (darker colors) vs. demographics only (lighter colors). The full-persona information only marginally helps improve individual-level accuracy. Moreover, digital twins are only as good as XGBoost trained on approximately 75 human responses on average, suggesting that it is harder for the LLM to leverage these data to make out-of-distribution predictions that would be accurate at the individual level.
In sum, this analysis suggests that the twins’ blueprint data, despite being based on a large number of scales and measures developed over decades of social science research, are far from sufficient to fully align digital twins’ answers to those of their human counterparts. Predicting human behavior in our context is probably intrinsically challenging and performance expectations need to be appropriately calibrated, and/or that new scales and measures may need to be developed to better capture human cognition. This analysis further suggests that the blueprint data is relatively more informative of variations across people and that LLMs are able to leverage these data to make out-of-distribution predictions that capture some variation across people (e.g., whether one person will score higher or lower than another), while not being necessarily more accurate at the individual level (e.g., whether one person would answer a 6 vs. a 7 on a scale).
Similarly, what role does the base LLM versus individual-level information play in shaping the behavior of digital twins and the distortions we observe? One may speculate, and test in future research, that several of the distortions we observe are due to properties of the base LLM being transferred to digital twins. For example, ideological biases and hyper-rationality likely stem from the base LLM powering the digital twins. It is similarly quite plausible that representation bias is linked to certain groups (e.g., higher education, moderate views) being better represented in the base model’s training data. And the insufficient individuation of answers by digital twins is probably the result of the homogeneity of the base model not being sufficiently offset by the individual-level information contained in digital twins.
Of course, a base model is required to enable digital twins to process and generate unstructured data, and to make out-of-distribution predictions without additional training data. But future research might find optimal ways to combine the capabilities of one (or potentially an ensemble of) base LLM(s) with rich individual-level information. We hope our data will encourage the development of improved pipelines for digital twins that address some of the limitations uncovered in our analysis.
@@
Finally, our findings also raise the question of how we should reasonably think about digital twins. In particular, based on our results, it may not be realistic to think about them as “clones” of humans, but rather as hyper-rational, quasi-omniscient versions of humans, with implicit values partly imbued by their base LLM. This perspective may influence which use cases are more promising for future research and practical applications. Perhaps use cases that consider digital twins as well-informed advisors are more promising than use cases that consider digital twins as carbon copies of their human counterparts.